Please help me with this! :)
The senior class at highschool A and highschool B planned seperate trips to the state fair. The senior class at highschool A rented and filled 4 vans and 3 busses with 125 students. Highschool B rented 4 vans and 8 busses with 240 students. Each van and each bus carried the same number of students. Find the number of students in each van and in each bus.
Thank you so much!

Answer by ankor@dixie-net.com(22740)

he senior class at highschool A and highschool B planned seperate trips to the state fair.
The senior class at highschool A rented and filled 4 vans and 3 busses with 125 students.
Highschool B rented 4 vans and 8 busses with 240 students.
Each van and each bus carried the same number of students. Find the number of students in each van and in each bus.
:
let v = no. of students in the van
Let b = no. in the bus
:
School A:
4v + 3b = 125
School B:
4v + 8b = 240
:
Subtract the 1st equation from the 2nd equation,
4v + 8b = 240
4v + 3b = 125
----------------subtraction eliminates v, find b
5b = 115
b = 115/5
b = 23 students in each bus
:
Find no. in the vans using the 1st equation
4v + 3(23) = 125
4v + 69 =
4v = 125 - 69
4v = 56
v = 56/4
v = 14 students in each van
:
Check solutions in 2nd equation
4(14) + 8(23) =
56 + 184 = 240 kids